OpenStudy (anonymous):

6a-(5+5a)=(3+a)-8 solve for variable A:all real numbers B:all intergers C:all real numbers

OpenStudy (anonymous):

What is difference b/w A and C? Anwer is A or C

OpenStudy (anonymous):

6a-(5+5a)=(3+a)-8 6a -5 - 5a = 3 + a - 8 a - 5 = a -5 -5 = -5 A: and C: all real numbers